OpenStudy (anonymous):

T(t)=63(.05)^(t/10)+19 . Where T is temperature, and t is time in minutes. How long did it take for a sandwich to reach an internal temputure of 30 degrees. I know that I have to substitute 30 into T. But I don't know how to solve for t. Please help and show steps.

OpenStudy (anonymous):

30 = 63(0.5)^(t/10) + 19 11 = 63(0.5)^(t/10) 11/63 = (0.5)6(t/10) ln(11/63) = (t/10)ln(0.5) Should be able to solve from there.

OpenStudy (anonymous):

I'm not able to solve from here. How would I do this. Please.

OpenStudy (anonymous):

I get where you got everything from what I"m having trouble with is isolating t.

OpenStudy (anonymous):

As of right now I have 11/63=(.5)^t/10

OpenStudy (anonymous):

Do you know logarithms?

OpenStudy (anonymous):

Or at least what they are?

OpenStudy (anonymous):

Nope. I think we learn that next year.

OpenStudy (jdoe0001):

makes me wonder why you're doing this now then :|

OpenStudy (anonymous):

Agreed, the only way to solve for the independent variable t is to use logarithms.

OpenStudy (anonymous):

Lol, Can you give me the jist of this? I probably have learned it but we don't call it logarithms.

OpenStudy (anonymous):

http://www.purplemath.com/modules/logs.htm

OpenStudy (jdoe0001):

so, you learned it "next year", but can't recall it in the present..... ooookkk

OpenStudy (anonymous):

I haven't learned it. Its in next years curriculum im told...

OpenStudy (anonymous):

We havent learned this. o.o Log business.

OpenStudy (jdoe0001):

so, this exercise doesn't apply then

OpenStudy (anonymous):

\[\log_{b}y = x ...... b^x=y \]

OpenStudy (anonymous):

Thanks for your help I guess.